How many atoms of hydrogen are in 2.46E+1 g of methane? Respond with the correct number of significant figures in scientific notation (Use E notation and only 1 digit before decimal e.g. 2.5E5 for 2.5 x 10⁵
Answer #1

Molar mass of CH4,

MM = 1*MM(C) + 4*MM(H)

= 1*12.01 + 4*1.008

= 16.042 g/mol

mass(CH4)= 24.6 g

use:

number of mol of CH4,

n = mass of CH4/molar mass of CH4

=(24.6 g)/(16.04 g/mol)

= 1.533 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 1.533 * 6.022*10^23 molecules

number of molecules = 9.235*10^23 molecules

1 molecule of CH4 has 4 atoms of Hydrogen

So,

number of H atoms = 4*number of CH4

= 4*9.235*10^23 atoms

= 3.69*10^24 atoms

Answer: 3.69E24 atoms